A helium-containing balloon has a volume of 10.6 L and a pressure of 1.22 atm at a temperature of 27.9C. What will the pressure inside the balloon be if the temperature is lowered to -11.3C? Assume that the gas behaves ideally. Express your answer in units of atmospheres (atm) using at least three significant figures.
Answer #1

SOLUTION :


As per Gas Law,


PV/T = R (gas constant)

If V remains unchanged, then,

P/T = constant


So,


P1 / T1 = P2 / T2

=> P2 = P1 * (T2 / T1) 

=> P2 = 1.22 ((273 - 11.3) / (273 + 27.9))

=> P2 = 1.061 atm 


So, pressure in balloon at decreased temperature of -11.3ºC  will be 1.061 atm (ANSWER).


answered by: Tulsiram Garg